OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2ð). cos 4x - cos 2x = 0 i dont get how to solve this

OpenStudy (lopezking1):

cos is a periodic function. the period is 2π, i.e. cosx=cos(x+2kπ) where k is an integer

OpenStudy (anonymous):

which integer do i use? 4 or 2

OpenStudy (lopezking1):

2pi/3 4pi/3 I think

OpenStudy (lopezking1):

Apply Cos C - Cos D formula cos C - Cos D = 2 sin(C+D)/2 Sin(D-C)/2

OpenStudy (anonymous):

what formula is rthat?!

OpenStudy (anonymous):

what is c and what is d\

OpenStudy (lopezking1):

lol it would be a long explanation but the answer would amount to either Sin3x=0 or Sin x=0 3x = n pi or x= n pi , where n =...-2,-1,0,1,2,3....

OpenStudy (lopezking1):

i hope this helps some how cos 4 x - cos 2 x = 0 2cos^2 2x-1 - cos 2x = 0 2u^2 -u -1 = 0 (u-1)(2u+1) = 0 u = 1; u = -1/2 cos 2x = 1 ; cos 2x = -1/2 2x = 0 ; 2x = pi-pi/3, 2x = pi + pi/3 x = 0 + npi; x = pi/3 + npi, x = 2pi/3 + npi x = 0, pi; x = pi/3, 4pi/3, x = 2pi/3, 5pi/3 so solutions will be : x = 0, pi/3, 2pi/3, 4pi/3, 5pi/3, pi

OpenStudy (lopezking1):

@lbouskila

OpenStudy (anonymous):

thank you!

OpenStudy (lopezking1):

:) welcome